# Optimal frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v1

@@ 1.3.1. Estimating the minimal condition number
We first estimate the minimal condition number of real $m\times 2$ matrices. Note that the condition number of a matrix $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{m}]^{T}\in\mathbb{R}^{m\times 2}$ does not change if the row vector $\bm{a}_{i}$ is replaced by $\bm{a}_{i}$ for any $i\in[m]:=\{1,\dotsc,m\}$. Hence, to estimate the minimal condition number, it is enough to consider the case when each $\bm{a}_{i}\in\mathcal{T}$.
@@
Assume that $\mathcal{A}=\{\bm{a}_{1},\ldots,\bm{a}_{m}\}\subset\mathcal{T}\subset{\mathbb{R}}^{2}$, where $\mathcal{T}$ is defined in (1.9). The set $\mathcal{A}$ is an optimal tight frame if and only if $\{f(\bm{a}_{1}),\dotsc,f(\bm{a}_{m})\}$ is the edge vectors set of an optimal $m$-gon $P\in\mathcal{P}_{m}$, where $f(\cdot)$ is defined in (1.10). In either case, we have
$\beta_{\bm{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}},$
where $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{m}]^{T}$ and $r(P)$ is defined in (1.7).
@@
    $\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}=\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
@@
    $\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}>\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
@@
    The harmonic frame $\bm{E}_{m}$ in (1.5) has the minimal condition number if and only if $m$ is odd. Moreover, $\bm{E}_{m}$ is the unique matrix with minimal condition number if and only if $m$ is a odd prime.
@@
The authors in [32] proved that the lower bound (1.3) is tight if $m\geq 3$ is an odd integer. Corollary 1.1 (i) generalizes this result by showing that the lower bound (1.3) is tight as long as $m\geq 3$ has an odd factor. Moreover, Corollary 1.1 (iii) shows that Conjecture 1.1 is false, i.e., the condition number of harmonic frame $\bm{E}_{m}$ is not minimal for each even integer $m\geq 4$.
@@
It is known that $P$ is the unique optimal $4$-gon, and we have $r(P)=\frac{1}{2+\sqrt{6}-\sqrt{2}}$ [29, 28]. Let $\bm{A}\in\mathbb{R}^{4\times 2}$ be defined as in (1.6). By Theorem 1.2 we see that the matrix $\bm{A}$ has the minimal condition number $\beta_{\bm{A}}^{\mathbb{R}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71$. According to (1.4), we have $\beta_{\bm{E}_{4}}^{\mathbb{R}}={\sqrt{2+\sqrt{2}}}\approx 1.84$, which is strictly greater than $\beta_{\bm{A}}^{\mathbb{R}}$. This is aligned with Corollary 1.1 (iii). The optimal convex $4$-gon $P$ and the rows of $\bm{A}$ are both illustrated in Figure 1.
@@ 1.3.2. Characterization of All Optimal Polygons
Let $m\geq 3$ be an integer. Define $\zeta_{m}:=e^{{\rm i}\pi/m}$. Determine all vectors $\bm{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ such that
$g(\bm{\varepsilon}):=\sum_{j=1}^{m}\varepsilon_{j}\zeta_{m}^{j}=0.$ (1.11)
@@
Let $m\geq 3$ be an integer with an odd factor. Set $\bm{\mu}_{j}:=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}$ for each $j\in[m]$. Then we have
@@
    If $\bm{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies (1.11), then there exists an optimal polygon $P\in{\mathcal{E}}_{m}$ with the edge vectors $\{\varepsilon_{1}\bm{\mu}_{1},\ldots,\varepsilon_{m}\bm{\mu}_{m}\}$. Such a polygon $P$ is unique up to translation.
@@
    Conversely, for any optimal polygon $P\in{\mathcal{E}}_{m}$, there exists a vector $\bm{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.11), such that the edge vectors of $P$ are $\{\varepsilon_{1}\bm{\mu}_{1},\ldots,\varepsilon_{m}\bm{\mu}_{m}\}$, after a proper scaling and rotation.
@@
For a vector $\bm{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$, we define the shifting operator $\mathcal{S}$ and the flip operator $\mathcal{R}$ as
$\mathcal{S}(\bm{\varepsilon})=(\varepsilon_{2},\ldots,\varepsilon_{m},-\varepsilon_{1})^{T}\quad\text{and}\quad\mathcal{R}(\bm{\varepsilon})=(\varepsilon_{m},\varepsilon_{m-1},\ldots,\varepsilon_{1})^{T}.$ (1.12)
For each $i\in\{1,2\}$, let $\bm{\varepsilon}_{i}\in\{\pm 1\}^{m}$ be a vector satisfying (1.11), and let $P_{i}\in{\mathcal{E}}_{m}$ be the corresponding optimal polygon with the edge set as described in Theorem 1.3 (i). A simple calculation shows that
@@
    if $P_{2}$ is obtained from $P_{1}$ through a proper translation and scaling, then $\bm{\varepsilon}_{2}=\bm{\varepsilon}_{1}$;
@@
    if $P_{2}$ is obtained from $P_{1}$ via a clockwise rotation of $\frac{k\pi}{m}$ for some integer $k\in[2m]$, then $\bm{\varepsilon}_{2}=\mathcal{S}^{k}(\bm{\varepsilon}_{1})$;
@@
    if $P_{2}$ is obtained from $P_{1}$ by a flip with respect to the $y$-axis, then $\bm{\varepsilon}_{2}=\mathcal{S}(\mathcal{R}(\bm{\varepsilon}_{1}))$.
Since we treat polygons as equivalent under scaling, translations, rotations, reflections, or any combination of these transformations, we consider two vectors $\bm{\varepsilon}_{1},\bm{\varepsilon}_{2}\in\{\pm 1\}^{m}$ to be equivalent if one can be obtained from the other by applying a finite sequence of the shifting operator $\mathcal{S}$ and the flip operator $\mathcal{R}$. Under this convention, the number of distinct solutions to Problem 1.1 equals that of Problem 1.2.
@@ 1.3.3. Characterization of All Optimal Frames
We next characterize all optimal frames with $m$ vectors when $m\geq 3$ has an odd factor. For convenience, for any vector ${\bm{\nu}}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$ with $t\geq 0$ and $\phi\in(0,\frac{\pi}{2}]$, we define ${\bm{\nu}}^{\perp}:=t\cdot(\cos(\phi+\frac{\pi}{2}),\sin(\phi+\frac{\pi}{2}))^{T}\in\mathcal{T}\subset{\mathbb{R}}^{2}$ .
@@
Assume that $m\geq 3$ is an integer with an odd factor. Let $\mathcal{T}_{m}$ denote the set of all optimal frames with $m$ vectors in $\mathcal{T}$, where $\mathcal{T}$ is defined in (1.9). Set ${\bm{\nu}}_{j}:=(\cos\frac{j\pi}{2m},\sin\frac{j\pi}{2m})^{T}$ for $j\in[m]$. Then we have
@@
    Assume that ${\mathcal{A}}=\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\in\mathcal{T}_{m}$. Then $\mathcal{A}$ forms a tight frame, and all vectors in $\mathcal{A}$} have the same norm, i.e.,
    $\|{\bm{a}}_{1}\|_{2}=\|{\bm{a}}_{2}\|_{2}=\cdots=\|{\bm{a}}_{m}\|_{2}.$
@@
    If $\bm{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies (1.11), then the frame ${\mathcal{A}}=\{{\bm{\nu}}_{j}:j\in I\}\cup\{{\bm{\nu}}_{j}^{\perp}:j\in[m]\backslash I\}$ is an optimal frame, i.e., $\mathcal{A}\in\mathcal{T}_{m}$, where $I=\{j\in[m]:\varepsilon_{j}=1\}$.
@@
    For any optimal frame ${\mathcal{A}}\in{\mathcal{S}}_{m}$, there exists a vector $\bm{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.11), such that ${\mathcal{A}}=\{{\bm{\nu}}_{j}:j\in I\}\cup\{{\bm{\nu}}_{j}^{\perp}:j\in[m]\backslash I\}$ where $I=\{j\in[m]:\varepsilon_{j}=1\}$, after a proper scaling and rotation.
@@
Theorem 1.4 establishes that any optimal frame ${\mathcal{A}}=\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset{\mathbb{R}}^{2}$ forms a tight frame when $m\geq 3$ has an odd factor. Inspired by this result, we propose the following conjecture: Any optimal frame ${\mathcal{A}}=\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset{\mathbb{R}}^{d}$ for phase retrieval in ${\mathbb{R}}^{d}$ must be a tight frame, provided $m\geq 2d-1$.
@@
Let $m\geq 3$ be an integer with an odd factor. A brute-force algorithm can be used to generate all optimal polygons in ${\mathcal{E}}_{m}$, up to scaling, translation, rotation, and reflection. It can also generate all optimal frames in $\mathcal{T}_{m}$, up to scaling, rotation, and reflection  [^1] The MATLAB and Maple implementations of the algorithm are available at https://github.com/zxynhy/Optimal-polygon-for-the-perimeter-maximizing-isodiametric-problem.git.. The algorithm begins by examining all vectors $\bm{\varepsilon}\in\{\pm 1\}^{m}$ that satisfy (1.11). Then Theorem 1.3 and Theorem 1.4 are applied. Take $m=12$ as an example. After checking all vectors $\bm{\varepsilon}\in\{\pm 1\}^{12}$ for which (1.11) holds, by Theorem 1.3, we identify all optimal polygons in ${\mathcal{E}}_{12}$ (up to scaling, translation, rotation and flip), shown in Figure 2 (a1) and (a2). Also, by Theorem 1.4, all the optimal frames in $\mathcal{T}_{12}$ (up to scaling, rotation and flip) are obtained, shown in Figure 2 (b1) and (b2). Denote by $\mathcal{T}_{m}$ the number of distinct frames in $\mathcal{T}_{m}$, modulo scaling, rotation and flip. Table 1 lists the number $\mathcal{T}_{m}$ for $3\leq m\leq 15$. In particular, the cases $\mathcal{T}_{4}=\#\mathcal{T}_{8}=1$ follow from established results in [29, 28, 4, 9]. A promising direction for future research is to design an efficient algorithm capable of finding all solutions to Problem 1.2.

### v2

@@ 1.3.1. Estimating the minimal condition number
We first estimate the minimal condition number of real $m\times 2$ matrices. Note that the condition number of a matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{R}^{m\times 2}$ does not change if the row vector $\boldsymbol{a}_{i}$ is replaced by $\boldsymbol{a}_{i}$ for any $i\in[m]:=\{1,\dotsc,m\}$. Hence, to estimate the minimal condition number, it is enough to consider the case when each $\boldsymbol{a}_{i}\in\mathcal{T}$.
@@
Assume that $\mathcal{A}=\{\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}\}\subset\mathcal{T}\subset{\mathbb{R}}^{2}$, where $\mathcal{T}$ is defined in (1.9). The set $\mathcal{A}$ is an optimal tight frame if and only if $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge vectors set of an optimal $m$-gon $P\in\mathcal{P}_{m}$, where $f(\cdot)$ is defined in (1.10). In either case, we have
$\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}},$
where $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}$ and $r(P)$ is defined in (1.7).
@@
    $\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
@@
    $\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}>\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
@@
    The harmonic frame $\boldsymbol{E}_{m}$ in (1.5) has the minimal condition number if and only if $m$ is odd. Moreover, $\boldsymbol{E}_{m}$ is the unique matrix with minimal condition number if and only if $m$ is a odd prime.
@@
The authors in [32] proved that the lower bound (1.3) is tight if $m\geq 3$ is an odd integer. Corollary 1.1 (i) generalizes this result by showing that the lower bound (1.3) is tight as long as $m\geq 3$ has an odd factor. Moreover, Corollary 1.1 (iii) shows that Conjecture 1.1 is false, i.e., the condition number of harmonic frame $\boldsymbol{E}_{m}$ is not minimal for each even integer $m\geq 4$.
@@
It is known that $P$ is the unique optimal $4$-gon, and we have $r(P)=\frac{1}{2+\sqrt{6}-\sqrt{2}}$ [29, 28]. Let $\boldsymbol{A}\in\mathbb{R}^{4\times 2}$ be defined as in (1.6). By Theorem 1.2 we see that the matrix $\boldsymbol{A}$ has the minimal condition number $\beta_{\boldsymbol{A}}^{\mathbb{R}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71$. According to (1.4), we have $\beta_{\boldsymbol{E}_{4}}^{\mathbb{R}}={\sqrt{2+\sqrt{2}}}\approx 1.84$, which is strictly greater than $\beta_{\boldsymbol{A}}^{\mathbb{R}}$. This is aligned with Corollary 1.1 (iii). The optimal convex $4$-gon $P$ and the rows of $\boldsymbol{A}$ are both illustrated in Figure 1.
@@ 1.3.2. Characterization of all optimal polygons
Let $m\geq 3$ be an integer. Define $\zeta_{m}:=e^{{\rm i}\pi/m}$. Determine all vectors $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ such that
$g(\boldsymbol{\varepsilon}):=\sum_{j=1}^{m}\varepsilon_{j}\zeta_{m}^{j}=0.$ (1.11)
@@
Let $m\geq 3$ be an integer with an odd factor. Set $\boldsymbol{\mu}_{j}:=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}$ for each $j\in[m]$. Then we have
@@
    If $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies (1.11), then there exists an optimal polygon $P\in{\mathcal{E}}_{m}$ with the edge vectors $\{\varepsilon_{1}\boldsymbol{\mu}_{1},\ldots,\varepsilon_{m}\boldsymbol{\mu}_{m}\}$. Such a polygon $P$ is unique up to translation.
@@
    Conversely, for any optimal polygon $P\in{\mathcal{E}}_{m}$, there exists a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.11), such that the edge vectors of $P$ are $\{\varepsilon_{1}\boldsymbol{\mu}_{1},\ldots,\varepsilon_{m}\boldsymbol{\mu}_{m}\}$, after a proper scaling and rotation.
@@
For a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$, we define the shifting operator $\mathcal{S}$ and the flip operator $\mathcal{R}$ as
$\mathcal{S}(\boldsymbol{\varepsilon})=(\varepsilon_{2},\ldots,\varepsilon_{m},-\varepsilon_{1})^{T}\quad\text{and}\quad\mathcal{R}(\boldsymbol{\varepsilon})=(\varepsilon_{m},\varepsilon_{m-1},\ldots,\varepsilon_{1})^{T}.$ (1.12)
For each $i\in\{1,2\}$, let $\boldsymbol{\varepsilon}_{i}\in\{\pm 1\}^{m}$ be a vector satisfying (1.11), and let $P_{i}\in{\mathcal{E}}_{m}$ be the corresponding optimal polygon with the edge set as described in Theorem 1.3 (i). A simple calculation shows that
@@
    if $P_{2}$ is obtained from $P_{1}$ through a proper translation and scaling, then $\boldsymbol{\varepsilon}_{2}=\boldsymbol{\varepsilon}_{1}$;
@@
    if $P_{2}$ is obtained from $P_{1}$ via a clockwise rotation of $\frac{k\pi}{m}$ for some integer $k\in[2m]$, then $\boldsymbol{\varepsilon}_{2}=\mathcal{S}^{k}(\boldsymbol{\varepsilon}_{1})$;
@@
    if $P_{2}$ is obtained from $P_{1}$ by a flip with respect to the $y$-axis, then $\boldsymbol{\varepsilon}_{2}=\mathcal{S}(\mathcal{R}(\boldsymbol{\varepsilon}_{1}))$.
Since we treat polygons as equivalent under scaling, translations, rotations, flips, or any combination of these transformations, we define two vectors $\boldsymbol{\varepsilon}_{1},\boldsymbol{\varepsilon}_{2}\in\{\pm 1\}^{m}$ to be equivalent if one can be obtained from the other by applying a finite sequence of the shifting operator $\mathcal{S}$ and the flip operator $\mathcal{R}$. A simple calculation shows that ${\mathcal{R}}^{2}={\mathcal{S}}^{2m}={\rm Id}$, where $\rm Id$} denotes the identity operator. Furthermore, we have ${\mathcal{S}}^{k}{\mathcal{R}}={\mathcal{R}}{\mathcal{S}}^{2m-k}$ for any integer $k\in[1,2m]$. Hence, $\boldsymbol{\varepsilon}_{1}$ is equivalent to $\boldsymbol{\varepsilon}_{2}$ if and only if ${\mathcal{R}}^{k_{1}}{\mathcal{S}}^{k_{2}}(\boldsymbol{\varepsilon}_{1})=\boldsymbol{\varepsilon}_{2}$ for some $k_{1}\in\{0,1\}$ and $k_{2}\in\{0,1,\ldots,2m-1\}$. Under this convention, the number of distinct solutions to Problem 1.1 equals that of Problem 1.2.
@@ 1.3.3. Characterization of all optimal frames
We next characterize all optimal frames with $m$ vectors when $m\geq 3$ has an odd factor. For convenience, for any vector ${\boldsymbol{\nu}}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$ with $t\geq 0$ and $\phi\in(0,\frac{\pi}{2}]$, we define ${\boldsymbol{\nu}}^{\perp}:=t\cdot(\cos(\phi+\frac{\pi}{2}),\sin(\phi+\frac{\pi}{2}))^{T}\in\mathcal{T}\subset{\mathbb{R}}^{2}$ .
@@
Assume that $m\geq 3$ is an integer with an odd factor. Let $\mathcal{T}_{m}$ denote the set of all optimal frames with $m$ vectors in $\mathcal{T}$, where $\mathcal{T}$ is defined in (1.9). Set ${\boldsymbol{\nu}}_{j}:=(\cos\frac{j\pi}{2m},\sin\frac{j\pi}{2m})^{T}$ for $j\in[m]$. Then we have
@@
    Assume that ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\in\mathcal{T}_{m}$. Then $\mathcal{A}$ forms a tight frame, and all vectors in $\mathcal{A}$} have the same norm, i.e.,
    $\|\boldsymbol{a}_{1}\|_{2}=\|\boldsymbol{a}_{2}\|_{2}=\cdots=\|\boldsymbol{a}_{m}\|_{2}.$
@@
    If $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies (1.11), then the frame ${\mathcal{A}}=\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$ is an optimal frame, i.e., $\mathcal{A}\in\mathcal{T}_{m}$, where $I=\{j\in[m]:\varepsilon_{j}=1\}$.
@@
    For any optimal frame ${\mathcal{A}}\in{\mathcal{S}}_{m}$, there exists a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.11), such that ${\mathcal{A}}=\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$ where $I=\{j\in[m]:\varepsilon_{j}=1\}$, after a proper scaling and rotation.
@@
Theorem 1.4 establishes that any optimal frame ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset{\mathbb{R}}^{2}$ forms a tight frame when $m\geq 3$ has an odd factor. Inspired by this result, we propose the following conjecture: Any optimal frame ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset{\mathbb{R}}^{d}$ for phase retrieval in ${\mathbb{R}}^{d}$ must be a tight frame, provided $m\geq 2d-1$.
@@
Let $m\geq 3$ be an integer with an odd factor. A brute-force algorithm can be used to generate all optimal polygons in ${\mathcal{E}}_{m}$, up to scaling, translation, rotation, and flip. It can also generate all optimal frames in $\mathcal{T}_{m}$, up to scaling, rotation, and flip  [^1] The MATLAB and Maple implementations of the algorithm are available at https://github.com/zxynhy/Optimal-polygon-for-the-perimeter-maximizing-isodiametric-problem.git.. The algorithm begins by examining all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{m}$ that satisfy (1.11). Then Theorem 1.3 and Theorem 1.4 are applied. Take $m=12$ as an example. After checking all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{12}$ for which (1.11) holds, by Theorem 1.3, we identify all optimal polygons in ${\mathcal{E}}_{12}$ (up to scaling, translation, rotation and flip), shown in Figure 2 (a1) and (a2). Also, by Theorem 1.4, all the optimal frames in $\mathcal{T}_{12}$ (up to scaling, rotation and flip) are obtained, shown in Figure 2 (b1) and (b2). Denote by $\mathcal{T}_{m}$ the number of distinct frames in $\mathcal{T}_{m}$, modulo scaling, rotation and flip. Table 1 lists the number $\mathcal{T}_{m}$ for $3\leq m\leq 15$. In particular, the cases $\mathcal{T}_{4}=\#\mathcal{T}_{8}=1$ follow from established results in [29, 28, 4, 9]. A promising direction for future research is to design an efficient algorithm capable of finding all solutions to Problem 1.2.